# Provable Emergence of Deep Neural Collapse and Low-Rank Bias in $L^{2}$-Regularized Nonlinear Networks

### v2

@@ Appendix B Global minima of representation cost for feedforward networks
Lemma B.1 (Optimal single layer for feedforward representation cost with singular value constraint).
@@
Assume without loss of generality that the singular values of both $A,B$ are ordered in a descending manner. Now, since $QQ^{\top}=I$, this is a classical form of Brockett’s cost function [4], whose solution is given by $Q^{*}=[I_{r},0_{r,m-r}]$ by using Theorem 2.1 in [23], which is a generalization of the classical Ky-Fan trace inequality. Then, we have
@@
Now, to prove the existence of the path $\gamma(t)$, we assume $n=m=N$ and, without loss of generality, we consider $C=1$. From Equation˜19 it is sufficient to find a path for the singular values $(s_{i}(t))$, and for the matrix $Q(t)\in O_{N}(\mathbb{R})$ such that they make $\mathscr{L}_{loc}$ decrease and reach the optimal value.
The existence of a descent path in $Q$, for fixed $\Sigma_{A},\Sigma_{B}^{+}$, is guaranteed by Lemma˜B.2. Then, we can assume w.l.o.g. that $Q=I$ by assuming that the singular values of $A$ and $B$ are both ascendingly ordered. So we concatenate the path in $Q$ with the following one for the singular values of $B$:
@@
The path in $Q$ decreases $\mathscr{L}$ by Lemma˜B.2. Additionally, by taking the derivative in $t$ of $\mathscr{L}_{loc}$ along the path (20), from (19) we have
@@
First, we observe that the structure of all stationary and in particular, of global minima and saddle points is studied in [3, Theorem 4]. From this last result all the minima are global minima, and all saddle points are strict under the hypothesis that all the eigenvalues of both $D_{1}$ and $D_{2}$ are distinct. Let us now consider the Riemannian gradient flow
@@
where $\mathrm{grad}$ is the gradient induced by the Frobenius inner product on $O_{N}(\mathbb{R})$. We now use [2, Theorem 6.3], which guarantees almost sure (with respect to the initial condition) convergence of the gradient flow to the local minima. Since we know that all local minima are global, we get the desired thesis.
@@
Theorem (Theorem˜4.2).
Let $\sigma_{2},\dots,\sigma_{L}$ be a set of entrywise homeomorphisms, $\sigma_{1}$ entrywise nonlinear with $\sigma_{l}(x)=0$ if and only if $x=0$. Assume $K=n_{L}\leq\dots\leq n_{1}=N$. Let $\mathcal{S}\subset\mathbb{R}^{n_{1}\times N}\times\dots\times\mathbb{R}^{n_{L}\times N}$ be as in Definition˜4.1. Then, the global minima of the optimization problem
$\min_{Z\in\mathcal{S}}\mathscr{L}(Z)$
$\bar{Z}^{l}\propto O_{l}\sigma_{l+1}^{-1}(\bar{Z}^{l+1}),\quad O_{l}\in\mathbb{R}^{n_{l}\times n_{l+1}},O_{l}^{\top}O_{l}=I,\quad\forall\,l=1,\dots,L-1.$
@@
$S_{l}(Z^{l-1},Z^{l+1})\subseteq\mathcal{S}_{loc}(\sigma_{l+1}^{-1}(Z^{l+1})).$ (21)
As a consequence we study the infimum in $Z^{1}$ and observe the following set of inequalities
@@
where we have used first (21), second, Lemma˜B.1, which proves the existence of the minimum in $\mathcal{S}_{loc}(\sigma_{2}^{-1}(Z^{2}))$ and characterizes its value, and third, condition (III) in Definition˜4.1, which guarantees the last inequality when applied to $\operatorname{rank}(\sigma_{2}^{-1}(Z^{2}))$. Next, we focus on the lower bound in (22) and define the following function, depending only on $Z^{2},\dots Z^{L}$:
@@
Thus, we proceed optimizing $Z^{2}$ in $\mathscr{L}_{1}$ as we did for $Z^{1}$ in $\mathscr{L}$ and obtain
@@
$\mathscr{L}_{i}(Z^{L},\dots,Z^{i}):=\frac{1}{2}\sum_{l=i+1}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}+\frac{i\operatorname{rank}(Y)}{2}.$ (24)
@@
Now we claim that there exists some $\bar{Z}\in\mathcal{S}$ such that $\mathscr{L}(\bar{Z})=(L/2)\operatorname{rank}(Y)$. Note that, as a consequence of the claim, we have that the minimum in $\mathcal{S}$ exists and that all the inequalities in (22) and (23) are equalities. In particular, because of Lemma˜B.1 any minimum $\bar{Z}$ has to satisfy
@@
for some sequence of matrices $O_{l}\in\mathbb{R}^{n_{l}\times n_{l+1}}$ such that $O_{l}^{\top}O_{l}=I$. We conclude by observing that $Y=Z^{L}$ satisfies DNC1 and so, applying (25) backwards from $L$ to $1$, also all the other layers $\bar{Z}^{l}$, at the minima, satisfy DNC1.
We are left to prove the claim. To this end, recall that $y_{j}=e_{i_{j}}$ for any $j$ and $\sigma_{l}(0)=0$ for any $l$. So, $\sigma_{L}^{-1}(y_{j})=\alpha_{L}e_{i_{j}}$ for any $j$ where $\alpha_{L}$ is a constant different from zero and independent of $j$. In particular we can consider $O_{L-1}^{T}=[\mathrm{Id},0]$ and $\bar{Z}^{L-1}=O_{L-1}\sigma_{L}^{-1}(Y)=\alpha[Y,0]^{T}$ and iteratively do the same for any $l$. The obtained $\bar{Z}$ is then included in $\mathcal{S}$ because each layer $Z^{l}$ and its counterimage $\sigma_{l}^{-1}(Z^{l})$ are a rescaled version of $Y$, therefore conditions (I),(II),(III) are satisfied. Moreover, by its definition, $\|\sigma_{l+1}^{-1}(\bar{Z}^{l+1})\bar{Z}^{l,+}\|^{2}=\operatorname{rank}(\sigma_{l+1}^{-1}(\bar{Z}^{l+1}))=\operatorname{rank}(Y)$, yielding $\mathscr{L}(\bar{Z})=(L/2)\operatorname{rank}(Y)$ and concluding the proof of the claim. ∎
B.1 Reachability of DNC configuration for almost every initial condition
A natural question arising from the last section is: from which initial configurations can the global minima be reached? In this section, we will prove not only that the $\operatorname{\mathrm{DNC1}}$ minima are global, but also that they are reachable from any interpolating initial condition in a related setting thanks to the result in Theorem˜4.3.
We are now ready for the main result about the reachability of rank-collapsed configurations through energy-decreasing paths:
Theorem (Theorem˜4.3).
Assume the intermediate widths in $\mathcal{S}$ satisfy the condition $K\leq n_{L}=n_{L-1}=\dots=n_{1}=N$ and that the $\sigma_{l}$’s are analytic. Let $Z(0)\in\mathcal{S}$ be a random initial interpolating condition, i.e., such that $Z^{L}(0)=Y$. Then, almost surely there exists a continuous path $\gamma(t)$ with $\gamma(0)=Z(0)$, $\gamma(1)$ global minimum of $\mathscr{L}$, and such that $\mathscr{L}\circ\gamma$ is continuous Lebesgue-almost everywhere (except for a finite set of times) and for all continuity times $s\geq t$ we have
$\mathscr{L}(\gamma(t))\geq\mathscr{L}(\gamma(s)).$
Moreover, all discontinuities $t_{1},\dots,t_{m}$ are removable and satisfy
$\underset{t\to t_{i}^{-}}{\lim}\mathscr{L}(\gamma(t))=\underset{t\to t_{i}^{+}}{\lim}\mathscr{L}(\gamma(t))>\mathscr{L}(\gamma(t_{i})),\quad\forall i=1,\dots,m.$
@@ B.1 Reachability of DNC configuration for almost every initial condition
Define therefore $\gamma_{1}(t)$ as in Lemma˜B.1, to move $Z^{1}(0)$ to $Z^{2}(0), OO$$^{\top}O=I$. We recall the notation $\mathscr{L}(Z)=\sum_{l=1}^{L-1}\mathscr{L}_{l}$, where $\mathscr{L}_{l}:=\frac{1}{2}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}$. For the second layer, concatenate this path with $\gamma_{2}^{2}(t)$ a path connecting $Z^{2}(0)$ with an optimum $Z^{2,*}(Z^{3})$ and $\gamma_{2}^{1}(t)\in Z^{1,*}(\gamma_{2}^{2}(t))=\gamma_{2}^{2}(t)$. For this second path, thanks to Lemma˜B.3 and the hypothesis that $\operatorname{rank}(\sigma_{2}^{-1}(Z^{2}(0)))=N$, we have that for all $t\in[0,1]$ but a finite set, $\operatorname{rank}(\sigma_{2}^{-1}(\gamma_{2}^{2}(t)))=N$, and therefore the function $\mathscr{L}_{2}(\gamma_{2}^{2}(t))+\operatorname{rank}(\sigma_{2}^{-1}(\gamma_{2}^{2}(t)))^{2}$ is continuous up to a finite number of removable discontinuities. In the discontinuities $t_{i}$, as shown in Lemma˜B.3, $\operatorname{rank}(\sigma_{2}^{-1}(\gamma_{2}^{2}(t_{i})))<N$, the jump is down. In the other points, we have
@@
where the last is because the rank of $\sigma_{2}^{-1}(\gamma_{2}^{2}(t))$ is constant along the path $\gamma_{2}^{2}(t)$. Again, inductively we can define a path $\gamma_{3}^{3}$ which makes $\mathscr{L}_{3}$ decrease and $\gamma_{3}^{2}(t)\in Z^{2,*}(\gamma_{3}^{3}(t)),\gamma_{3}^{1}(t)=Z^{1,*}(\gamma^{2}_{3}(t))=\gamma^{2}_{3}(t)$. Along this path, again, up to a finite number of times in which there is a jump down, we have
$\frac{d}{dt}\mathscr{L}(Z^{L}(0),\dots,\gamma_{3}^{3}(t),\gamma_{2}^{3}(t),\gamma^{1}_{3}(t))=\frac{d}{dt}\mathscr{L}_{3}(\gamma_{3}^{3}(t))\leq 0.$
At layer $l$, we generally have paths $\gamma_{l}^{l^{\prime}}(t)$ for which the required decrease condition is satisfied. By concatenating all of them we have the required path $\gamma$. At the last path, the rank collapses, at $t=1$. For $t\in[0,1]$ we have (again thanks to Lemma˜B.3 up to a finite number of points $\{t_{1},\dots,t_{k}\}$) that
$\frac{d}{dt}\mathscr{L}(Y,\gamma_{L-1}^{L-1}(t),\dots,\gamma_{L-1}^{1}(t))=$ $\frac{d}{dt}\mathscr{L}_{L-1}(\gamma_{L-1}^{L-1}(t))$ (26)
$+\sum_{l=1}^{L-2}\frac{d}{dt}\operatorname{rank}(\sigma_{l}^{-1}(\gamma_{L-1}^{l}(t)))\leq 0,\,\,\,\text{for\, a.e.}\,t\in[0,1).$
At $t=1$ instead, the rank collapses for all layers to $K$, therefore again decreasing the objective. ∎
@@
As a corollary of Theorem˜4.3 we have the following:
Corollary B.4 (No loss barrier for any finite stepsize).
In the setting of Theorem˜4.3, for Lebesgue-almost every $Z(0)\in\mathcal{S}$ there exists a sequence of curves $\gamma_{k}:[0,1]\to\mathcal{S}$ uniformly converging to the continuous $\gamma$ of Theorem˜4.3 such that the following hold for all $k\geq 1$:
  • $\gamma_{k}$ is piecewise constant in a set of disjoint intervals intervals $\{I_{j,k}:=[T_{j,k},T_{j+1,k})\}_{j=0,\dots,k-1}$ whose union is $[0,1]$, and with $0<T_{j+1,k}-T_{j,k}<\frac{1}{k}$ for all $j=0,\dots,k-1$;
  • $\gamma_{k}(0)=Z(0)$;
  • $\gamma_{k}(1)\in\underset{Z\in\mathcal{S}}{\arg\min}\,\mathscr{L}(Z)$;
  • $\mathscr{L}$ is decreasing, i.e.,
    $\mathscr{L}(\gamma_{k}(0))\geq\mathscr{L}(\gamma_{k}(T_{1,k})))\geq\mathscr{L}(\gamma_{k}(T_{2,k})))\geq\dots\geq\mathscr{L}(\gamma_{k}(T_{k,k})))\geq\mathscr{L}(\gamma_{k}(1)).$
@@
Consider the curve $\gamma$ from Theorem˜4.3 together with its finite number of discontinuities $E:=\{t_{i}\}_{i}$ in which
@@
Remark B.5 (Intuition on Corollary˜B.4).
We highlight that an interpretation of Corollary˜B.4, is in terms of numerical schemes. In particular, since $\gamma_{k}:[0,1]\to\mathcal{S}$ is piecewise constant, we can think of it as an ordered sequence of $k$ points $\Gamma_{k}:=(\gamma_{k}(0),\gamma_{k}(T_{1,k}),\dots,\gamma_{k}(T_{k-1,k}),\gamma_{k}(1))$, where $\gamma_{k}(0)=Z(0)$ and $\gamma_{k}(1)\in\underset{Z\in\mathcal{S}}{\arg\min}\,\mathscr{L}$. We know from Corollary˜B.4 that $0<T_{j+1,k}-T_{j,k}:=\Delta_{j,k}<\frac{1}{k}$, and we can think of $\Delta_{j,k}$ as the stepsize of an optimization scheme. With this heuristic intuition in mind, the result proven in the corollary essentially states that for Lebesgue-almost every initialization $Z(0)\in\mathcal{S}$, there exists a numerical scheme with arbitrarily small maximal stepsize such that $\mathscr{L}$ decreases along it. Notice that, if there was any loss barrier between $Z(0)$ the the global minimizers of $\mathscr{L}$ on $\mathcal{S}$, then this would not be possible: in particular, if there was a barrier between initialization and minimizers, then there would exist a small enough stepsize such that $\mathscr{L}$ would need to increase along the sequence.
B.2 Stability of the $\operatorname{\mathrm{DNC1}}$ configurations, proof of Theorem˜4.4
@@
Remark B.6.
Notice that of interpretation of Theorem˜4.4 is the implicit bias of fully connected networks in the regime $\|Y\|\ll\|X\|$, i.e. the regime in which $\lambda\to+\infty$.
A second interpretation can be formalized as follows (and proved in the exact same way): consider the representation cost
$\mathscr{L}^{[l_{1},l_{2}]}_{L}(Z):=\begin{cases}+\infty,\quad\text{if}\,\,Z^{L}\neq Y\\ \frac{1}{L}\sum_{l=l_{1}}^{l_{2}}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2},\quad\text{otherwise}\end{cases}$
and consider the difference (without loss of generality, for $Z^{L}=Y$)
$\mathscr{L}^{[0,L-1]}_{L}-\mathscr{L}^{[l^{*},L-1]}_{L}=\mathscr{L}^{[0,l^{*}-1]}_{L}=\frac{1}{L}\sum_{l=0}^{l^{*}-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}\underset{\Gamma}{\to}0,\quad L\to+\infty$
for any finite $l^{*}$, which $\Gamma$-converges to zero for $L\to\infty$. Therefore, in the regime of large depth, even in the absence of the piramidal architecture on the backbone, one expects the minimizers to satisfy exactly $\operatorname{\mathrm{DNC1}}$ from the first layer $l^{*}$ for which the set of representable features $Z^{l^{*}}$ contains the minimizer predicted in Theorem˜4.2 (or alternatively, as a special case, when the set of representable matrices $Z^{l^{*}}$ is the whole $\mathbb{R}^{n_{l^{*}}\times N}$). This observation formalizes the theory in more general cases, essentially stating that for deep enough architectures, the structure of the minimizers is close to the one predicted in Theorem˜4.2, where $Z^{1}$ is substituted with the minimal $l^{*}$ such that $Z^{l^{*}}$ is “representative enough”.

### v3

@@ Appendix B Global Minima of Representation Cost for Feedforward Networks
Lemma B.1 (Optimal single layer for feedforward representation cost).
@@
Assume without loss of generality that the singular values of both $A,B$ are ordered in a descending manner. Now, since $QQ^{\top}=I$, this is a classical form of Brockett’s cost function (Brockett, 1991), whose solution is given by $Q^{*}=[I_{r},0_{r,m-r}]$ by using Theorem 2.1 in (Liang et al., 2023), which is a generalization of the classical Ky-Fan trace inequality. Then, we have
@@
Now, to prove the existence of the path $\gamma(t)$, we assume $n=m=N$ and, without loss of generality, we consider $C=1$. From Equation˜16 it is sufficient to find a path for the singular values $(s_{i}(t))$, and for the matrix $Q(t)\in O_{N}(\mathbb{R})$ such that they make $\mathscr{L}_{loc}$ decrease and reach the optimal value.
The existence of a descent path in $Q$, for fixed $\Sigma_{A},\Sigma_{B}^{+}$, is guaranteed by Lemma˜B.4. Then, we can assume w.l.o.g. that $Q=I$ by assuming that the singular values of $A$ and $B$ are both ascendingly ordered. So we concatenate the path in $Q$ with the following one for the singular values of $B$:
@@
The path in $Q$ decreases $\mathscr{L}$ by Lemma˜B.4. Additionally, by taking the derivative in $t$ of $\mathscr{L}_{loc}$ along the path equation 17, from equation 16 we have
@@
Remark B.2 (Local lemma for backbone).
In the case of a deep-architecture, the second layer requires a dedicated analysis because of the double representation constraint with the first layer
$\min_{B\in\mathcal{S}_{loc}(A),\,\sigma_{2}^{-1}(B)Z^{1,+}Z^{1}=\sigma_{2}^{-1}(B)}\mathscr{L}_{loc}(B).$
In particular, if the columns of $Z^{1}$ are linearly independent, then the second representation constraint is trivially satisfied. In fact, in that case $Z^{1,+}Z^{1}=Z^{1,-1}Z^{1}=I_{N}$, reducing to the case presented in Lemma˜B.1. Thanks to this observation and Lemma˜B.3, the following analysis holds in the pyramidal architecture hypothesis and with a fixed $W_{1}$ which satisfies the assumptions of Lemma˜B.3.
Lemma B.3 (Generic linear independence after a wide layer).
Consider an analytic non-polynomial activation function $\sigma:\mathbb{R}\to\mathbb{R}$ and a set of data points $\{x_{i}\}_{i=1}^{N}\subset\mathbb{R}^{n}$ with $n\leq N\leq 2^{n-1}$. Then for, Lebesgue-almost every couple $(X,W)\in\mathbb{R}^{n\times N}\times\mathbb{R}^{N\times n}$ it holds that $\det(\sigma(WX))\neq 0$, where $X=[x_{1},\dots,x_{N}]\in\mathbb{R}^{n\times N}$ and $\sigma$ applies entrywise.
Proof.
Letting $Z(W,X):=WX$, consider the determinant
$\det(\sigma(WX))=:F(Z(W,X)),$
which is analytic in $(W,X)$. Therefore, if we can prove that $F$ is not identically zero, then its zero set in $\mathbb{R}^{N\times n}\times\mathbb{R}^{n\times N}$ has Lebesgue measure zero (Mityagin, 2020, Proposition 1), and we can conclude. To show this, we need to exhibit a single matrix $Z$ with $\mathrm{rank}(Z)\leq n$ for which $F(Z)\neq 0$. Following (Chu et al., 2024, Theorem 2.5), whenever $\sigma$ is non-polynomial, then $\sigma$ is not rank preserving ( see (Chu et al., 2024, Definition 2.2)), meaning that there exists a matrix $Z$ with $\operatorname{rank}(Z)\leq\log_{2}(N)+1\leq n$ such that $\operatorname{rank}(\sigma(Z))=N$. This concludes the proof. ∎
@@
First, we observe that the structure of all stationary and in particular, of global minima and saddle points is studied in (Brockett, 1989, Theorem 4). From this last result all the minima are global minima, and all saddle points are strict under the hypothesis that all the eigenvalues of both $D_{1}$ and $D_{2}$ are distinct. Let us now consider the Riemannian gradient flow
@@
where $\mathrm{grad}$ is the gradient induced by the Frobenius inner product on $O_{N}(\mathbb{R})$. We now use (Bah et al., 2022, Theorem 6.3), which guarantees almost sure (with respect to the initial condition) convergence of the gradient flow to the local minima. Since we know that all local minima are global, we get the desired thesis.
@@
B.1 Proof of Theorem˜4.2
@@
$\mathcal{S}_{l}(Z^{l-1},Z^{l+1})\subseteq\mathcal{S}_{loc}(\sigma_{l+1}^{-1}(Z^{l+1})),$ (18)
where we refer to equation 15 for the definition of $\mathcal{S}_{loc}(\sigma_{l+1}^{-1}(Z^{l+1}))$. As a consequence we study the infimum in $Z^{1}$ and observe the following set of inequalities
@@
where we have used first equation 18, second, Lemma˜B.1, which proves the existence of the minimum in $\mathcal{S}_{loc}(\sigma_{2}^{-1}(Z^{2}))$ and characterizes its value, and third, condition (III) in Definition˜4.1, which guarantees the last inequality when applied to $\operatorname{rank}(\sigma_{2}^{-1}(Z^{2}))$. Next, we focus on the lower bound in equation 19 and define the following function, depending only on $Z^{2},\dots Z^{L}$:
@@
Thus, we proceed optimizing $\mathscr{L}_{1}$ with respect to $Z^{2}$ as we did for $\mathscr{L}$ with respect to $Z^{1}$ and obtain
@@
$\mathscr{L}_{i}(Z^{L},\dots,Z^{i}):=\frac{1}{2}\sum_{l=i+1}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}+\frac{i\operatorname{rank}(Y)}{2}.$
@@
Now we claim that there exists some $\bar{Z}\in\mathcal{S}$ such that $\mathscr{L}(\bar{Z})=(L/2)\operatorname{rank}(Y)$. Note that, as a consequence of the claim, we have that the minimum in $\mathcal{S}$ exists and that all the inequalities in equation 19 and equation 20 are equalities. In particular, because of Lemma˜B.1 any minimum $\bar{Z}$ has to satisfy
@@
for some sequence of matrices $O_{l}\in\mathbb{R}^{n_{l}\times n_{l+1}}$ such that $O_{l}^{\top}O_{l}=I$. We conclude by observing that $Y=Z^{L}$ satisfies DNC1 and so, applying equation 21 backwards from $L$ to $1$, also all the other layers $\bar{Z}^{l}$, at the minima, satisfy DNC1.
We are left to prove the claim. To this end, recall that $y_{j}=e_{i_{j}}$ for any $j$ and $\sigma_{l}(0)=0$ for any $l$. So, $\sigma_{L}^{-1}(y_{j})=\alpha_{L}e_{i_{j}}$ for any $j$ where $\alpha_{L}$ is a constant different from zero and independent of $j$. In particular we can consider $O_{L-1}^{T}=[\mathrm{Id},0]$ and $\bar{Z}^{L-1}=O_{L-1}\sigma_{L}^{-1}(Y)=\alpha[Y,0]^{T}$ and iteratively do the same for any $l$. The obtained $\bar{Z}$ is then included in $\mathcal{S}$ because each layer $Z^{l}$, as its counterimage $\sigma_{l}^{-1}(Z^{l})$, are a rescaled version of $Y$. In particular, conditions (I),(II),(III) are satisfied. Moreover, by its definition, $\|\sigma_{l+1}^{-1}(\bar{Z}^{l+1})\bar{Z}^{l,+}\|^{2}=\operatorname{rank}(\sigma_{l+1}^{-1}(\bar{Z}^{l+1}))=\operatorname{rank}(Y)$, yielding $\mathscr{L}(\bar{Z})=(L/2)\operatorname{rank}(Y)$ and concluding the proof of the claim. ∎
B.2 Reachability of DNC configuration for almost every initial condition
A natural question arising from the last section is: from which initial configurations can the global minima be reached? In this section, we will prove not only that the $\operatorname{\mathrm{DNC1}}$ minima are global, but also that they are reachable from any interpolating initial condition in a related setting thanks to the result in Theorem˜4.4.
We are now ready for the main result about the reachability of rank-collapsed configurations through energy-decreasing paths.
B.3 Proof of Theorem˜4.4
@@ B.2 Reachability of DNC configuration for almost every initial condition
We recall the notation $\mathscr{L}(Z)=\sum_{l=1}^{L-1}\mathscr{L}_{l}$, where $\mathscr{L}_{l}:=\frac{1}{2}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}$. First observe that, from Lemma˜B.1, there exists a path $\gamma_{1}(t)$ moving $Z^{1}(0)$ to an optimal $Z^{1,*}(Z^{2}(0))=OZ^{2}(0)$, for some $O^{\top}O=I$. After having optimized the first layer, we can optimize the second one. To this end, we consider a path $\gamma_{2}^{2}(t)$ moving $Z^{2}(0)$ to an optimum $Z^{2,*}(Z^{3}(0))$, which in turn induces a path $\gamma_{2}^{1}(t)\in Z^{1,*}(Z^{2}(t))$ on the first layer, which keeps it optimal. For this second path $(\gamma_{2}^{1}(t),\gamma_{2}^{2}(t))$, thanks to Lemma˜B.5 and the hypothesis that $\operatorname{rank}(\sigma_{2}^{-1}(Z^{2}(0)))=N$, we have that for all $t\in[0,1]$ but a finite set, $\operatorname{rank}(\sigma_{2}^{-1}(\gamma_{2}^{2}(t)))=N$, and therefore the function $\mathscr{L}_{2}(\gamma_{2}^{2}(t))+\operatorname{rank}(\sigma_{2}^{-1}(\gamma_{2}^{2}(t)))^{2}$ is continuous up to a finite number of removable discontinuities. In the discontinuities $t_{i}$, as shown in Lemma˜B.5, $\operatorname{rank}(\sigma_{2}^{-1}(\gamma_{2}^{2}(t_{i})))<N$, the jump is down. In the other points, we have
@@
where the last is because the rank of $\sigma_{2}^{-1}(\gamma_{2}^{2}(t))$ is constant along the path $\gamma_{2}^{2}(t)$.
Inductively, after having optimized the first l1l-1 layers, we can define paths $\gamma_{l}^{l^{\prime}}(t)$ for all $l^{\prime}\leq l$ that move $Z^{l}(0)$ to an optimum while keeping all the previous layers optimal and satisfying the required decrease condition
$\frac{d}{dt}\mathscr{L}(Z^{L}(0),\dots,\gamma_{l}^{l}(t),\dots,\gamma_{l}^{2}(t),\gamma_{l}^{1}(t))=\frac{d}{dt}\mathscr{L}_{l}(\gamma_{l}^{l}(t))\leq 0.$
By concatenating all of the paths that we have considered we obtain the required path $\gamma$. Note that for the last path, the rank of all layers stays constant for almost all times, while collapsing to $K$ at $t=1$, therefore again decreasing the objective. ∎
@@
As a corollary of Theorem˜4.4 we have the following:
B.4 Proof of Corollary˜4.5
@@
Consider the curve $\gamma$ from Theorem˜4.4 together with its finite number of discontinuities $E:=\{t_{i}\}_{i}$ in which
@@
Remark B.6 (Intuition on Corollary˜4.5).
We highlight that an interpretation of Corollary˜4.5, is in terms of numerical schemes. In particular, since $\gamma_{k}:[0,1]\to\mathcal{S}$ is piecewise constant, we can think of it as an ordered sequence of $k$ points $\Gamma_{k}:=(\gamma_{k}(0),\gamma_{k}(T_{1,k}),\dots,\gamma_{k}(T_{k-1,k}),\gamma_{k}(1))$, where $\gamma_{k}(0)=Z(0)$ and $\gamma_{k}(1)\in\underset{Z\in\mathcal{S}}{\arg\min}\,\mathscr{L}$. We know from Corollary˜4.5 that $0<T_{j+1,k}-T_{j,k}:=\Delta_{j,k}<\frac{1}{k}$, and we can think of $\Delta_{j,k}$ as the stepsize of an optimization scheme. With this heuristic intuition in mind, the result proven in the corollary essentially states that for Lebesgue-almost every initialization $Z(0)\in\mathcal{S}$, there exists a numerical scheme with arbitrarily small maximal stepsize such that $\mathscr{L}$ decreases along it. Notice that, if there was any loss barrier between $Z(0)$ the the global minimizers of $\mathscr{L}$ on $\mathcal{S}$, then this would not be possible: in particular, if there was a barrier between initialization and minimizers, then there would exist a small enough stepsize such that $\mathscr{L}$ would need to increase along the sequence.
B.5 Stability of the $\operatorname{\mathrm{DNC1}}$ Configurations and Proof of Theorem˜4.6